OpenStudy (anonymous):

simplify: (2ab^2)(-5a^(-4)b^5

OpenStudy (callisto):

(2ab^2)(-5a^(-4)b^5 \[= (2)(-5)a^{1-4}b^{2+5}\] = ..? Note: \(a^{-n} = \frac{1}{a^n}\)

OpenStudy (anonymous):

-10a1/4b^7???

OpenStudy (callisto):

Nope, not really. What is 1-4?

OpenStudy (anonymous):

-3

OpenStudy (callisto):

So, \(=-10a^{-3}b^7\) You need to know that \(a^{-n} = \frac{1}{a^n}\) By this, we can get \(=\frac{-10b^7}{a^{3}}\) Got it?

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

thank youu